OpenStudy (anonymous):

Find the length of the loop of the given curve. x = 3t - t^3 y = 3t^22 I know the formula, I just dont remember how to find the limits for t.

OpenStudy (anonymous):

look in the alternate forms tab

OpenStudy (anonymous):

that gives me zeros but not the closed interval. Am I missing something obvious?

OpenStudy (jamesj):

if this is a closed look, then we know that at t = a and b, (x,y)(t) is the same. So in particular, give the form of y(t), it must be that a = -b Now given that, is it the case that x(a) = x(b)? I.e., x(a) = x(-a) => 3a - a^3 = -3a + a^3 => 2a^3 - 6a = 0 => 2a (a^2 - 3) = 0 Hence the loop closes over the interval in t \[[-\sqrt{3}, \sqrt{3}]\]

OpenStudy (jamesj):

(Note that a = 0 is not a solution for us because that gives us no pair of points in the y-ordinate function.)